OpenStudy (anonymous):

Find the indicated value of the function. f(x) = √ (x+8) - x - 1 ; f(√ (2)+1) = ?

OpenStudy (anonymous):

is it \[f(x)=\sqrt{x+8}-x-1\]?

OpenStudy (anonymous):

No, that didn't work either.

OpenStudy (anonymous):

i was not saying that was the answer, i was asking if that was the function

OpenStudy (anonymous):

I'm not sure how to do the problem at all. I tried plugging in sqrt{2}+1 for each x.

OpenStudy (anonymous):

that is exactly what you need to do

OpenStudy (anonymous):

When I do that it doesn't say my answer is right.

OpenStudy (anonymous):

\[\sqrt{\sqrt{2}+1+8}-(\sqrt{2}+1)-1\]

OpenStudy (anonymous):

don't forget two distribute the minus sign and combine like terms

OpenStudy (anonymous):

Do you need to do everything under the radical first then distribute?

OpenStudy (anonymous):

there is not much to do under the radical besides saying \(1+8=9\)

OpenStudy (anonymous):

\[\sqrt{\sqrt{2}}+9 - \sqrt{2}\]

OpenStudy (anonymous):

Is that the answer?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

you need the distributive property for \(-(\sqrt{x}+1)\)

OpenStudy (anonymous):

remove parentheses, get \(-\sqrt{x}-1\) then subtract the 1 in the function and get \[-\sqrt{2}-2\] as the second part

OpenStudy (anonymous):

Where do you put that?

OpenStudy (anonymous):

\[\sqrt{\sqrt{2}+9}-\sqrt{2}-2\]Would it look like this